Find the equation of the perpendicular bisector of the line segment joining the points (3,4) and (-1.2).


 

Grade:upto college level

1 Answers

Sunil Kumar
9 years ago

Perpendicualr bisector means the line will have -ve reciprocal slope wrt the given line and it will pass through the mid-point of the line formed by the given coordinates.

So , 

slope of the line formed by coordinates of (3,4) and (-1/2) , m1 = (2-4)/(-1-3)=1/2

slope of the perpendicular bisector , m2= -2

Mid-point coordinates = (1,3)

Thus the equation of the line : y-3=-2(x-1) 

  or, 2x+y-5=0 is the equation of the line.

 

Thanks & Regards
Sunil Kumar, 
askIITians Faculty
M.Sc., IIT Kharagpur